# Same or Not? Enhancing Visual Perception in Vision-Language Models

### v1

@@ 5.3 Additional Results
To ensure post-training on Twin does not erode general visual reasoning, we report results on popular VQA benchmarks in Sec. 5.2. These benchmarks span common sense understanding (Seed [seedbench], Mmmu [mmmu], Nlvr2 [nlvr2]), spatial reasoning and grounding (Realworld [realworldqa], Cv-Bench [cambrian1], Blink [blink]), hallucination and bias (Pope [pope], VlmBias [vlmbias]), OCR (TextVQA [textvqa]), and chart/diagram understanding (ChartQA [chartqa], Ai2d [ai2d]). Although not our main focus, models post-trained on Twin retain performance on general VQA benchmarks, with small improvements on most for both Qwen2.5-VL ($1.4\%$ on Nlvr2; $0.8\%$ on Ai2d) and InternVL3.5 ($0.4\%$ on Realworld; $0.4\%$ on Cv-Bench).
Mmlu [mmlu] HellaSwag [hellaswag] Gsm8K [gsm8k]
Qwen2.5 VL 3B [qwen2.5vl] 64.7 66.4 70.8
@@
Beyond vision, Tab. 6 reports performance on popular text-only benchmarks covering different skills: Mmlu [mmlu] for common knowledge, HellaSwag [hellaswag] for sentence completion, and Gsm8K [gsm8k] for mathematical reasoning. We find that post-training leaves performance largely unchanged on these benchmarks: $64.7\%$ vs $64.6\%$ for Qwen2.5-VL on Mmlu; $57.8\%$ vs $58.1\%$ for InternVL3.5 on Gsm8K. These results reaffirm prior findings that RL-based tuning preserves model skills [rlsrazor, sftmemorizesrlgeneralizes].
@@ A.1 Evaluation Benchmarks
ILIAS [ilias] is a large-scale test dataset of instance-level image retrieval. It predominantly features images of retail products taken in various contexts, backgrounds, and lighting. For both pair and multi queries, we source images from the core_db split, using the __key__ field to determine instance identity.
Google Landmarks v2 (Landmarks) [landmarksv2] is a landmark recognition dataset featuring human-made and natural landmarks. The dataset has been used in both classification and retrieval settings. To ensure sufficient images per landmark to for multi queries, we source images from the train split and use the label field to determine landmark identity.
MET [met] is an image retrieval dataset featuring artwork from the Metropolitan Museum of Art in New York. The dataset features images of the same art piece or sculpture from varying viewpoints, emphasizing multi-view consistency in retrieval. We use the mini_met set to construct both pair and multi queries.
CUB [cub] is a fine-grained classification dataset that focuses on identifying bird species from images. We use the test split for all queries and create pair and multi queries based on the specie of bird.
Inquire [inquire] is a benchmark for natural world image retrieval, featuring images of animal and plant species sourced from iNaturalist [inaturalist]. We source images from the validation split and use the inat24_species_name field to determine instance identity for both pair and multi queries.
@@ A.2 Fgvqa Prompts
We include all prompts used on all datasets for both pair and multi queries in Figs. 12 to 17. All prompts are identical in format, except for the domain-specific instance names (e.g. bird/object/artwork/species). Additionally, for benchmarks featuring object instances (Twin-Eval, Ilias [ilias]), we found it necessary to provide the definition of an object instance as models were incorrectly labeling different objects of the same-category (e.g. two earbuds of different colors) as the same.
@@ B.1 Instance Sourcing
We used the Amazon Reviews dataset [amazon-reviews] to extract a large set of product listings, which would serve as candidate objects, and associated noisy review images. We categorized these objects into supercategories based on their associated category in the Amazon store (e.g. “Electronics”) and further into categories based on the product title (e.g. “Speaker”). We aimed to maximize diversity in our instances and thus selected 36 categories to draw from. We subsequently tasked human annotators with removing products that either lacked visual consistency (e.g. the same product being sold in multiple colors) or did not adhere to the object category (e.g. “earbud replacement tips” instead of “earbuds”).
Figure 8: Additional Synthetic Negatives. For each pair, we show a real image of the object instance and a synthetic image of the same instance generated with Dreambooth [dreambooth]. We find the synthetic negatives capture the “gist” of the instance, but are not faithful to fine details such as color, geometry, and texture – making them a strong additional set of hard negative pairs.
@@ B.3 Hard Negative Pair Assignment
Once we have collected a set of object instances, we construct hard negative pairs – pairs of distinct objects that appear similar. We collect these hard negatives with the help of human annotators. First, using CLIP [clip], for each object instance, we compute pairwise cosine similarities between the image embeddings of that instance and all other instances. We use these similarities to shortlist $k$ visually similar candidates for each instance, with $k$ being a random number between $1-2\times$ the number of positive images of that instance. Human annotators then select the final pairs, omitting pairs they deem to be too easy.
@@ B.4 Additional Details on Generated Negatives
We augment the set of negative pairs with additional synthetically generated images. Augmenting with generated negatives allows us to produce additional negative pairs without extra data collection. We follow the procedure outlined in Dreambooth [dreambooth]. For each object instance, we sample 3 to 5 images for each instance to serve as grounding images. We set the initializer tokens to the generic category associated with the instance (e.g. “vase”, “headphones”). Our prompts to Dreambooth are: [“an image of {category}”, “similar to {category}”, “a picture of{category}, “show me a {category}, “here is a {category}”] replacing {category} with the category of the instance. Finally, we task human annotators with validating the generated samples. Annotators are asked to remove generations that either do not appear similar to the instance or are indistinguishable from the instance – yielding a set of hard negative pairs.
@@ Appendix C Additional Training Details
We include additional details for all model training. In Sec. C.1, we provide an overview of GRPO [deepseekmath], our main post-training method, and our training setup. We provide additional details on supervised fine-tuning (Sec. 5.2) in Sec. C.2. We include all hyperparameters in Tab. 8 and Tab. 9.
@@ C.1 GRPO
To post-train VLMs on Twin, we use Group Relative Policy Optimization (GRPO; [deepseekmath]). We provide an overview of this method below.
@@
Optimization. We optimize our VLM $\pi_{\theta}$ via GRPO [deepseekmath]. For each image pair $(I_{1},I_{2})$ with ground truth label $y$, we sample $G$ responses: $\mathcal{O}=\{(I_{1},I_{2},r^{(i)},\hat{y}^{(i)})\}^{G}_{i=1}$ and compute centralized advantages:
@@
Implementation. We train Qwen2.5-VL-3B-Instruct [qwen2.5vl] and InternVL3.5-1B-Instruct [internvl3] on Twin, chosen as leading open-source VLMs at the 3B and 1B scales. We do not freeze any part of the model, and train on $4$ A100 GPUs for 1 epoch. We use a batch size of $480$, group size $5$ and learning rate $10^{-6}$. We build on the verl [verl] repository for training. We detail all hyperparameters in Tab. 8.
@@ C.2 Supervised Fine-Tuning
We compare post-training methods in Sec. 5.2. We provide additional details for supervised fine-tuning (SFT) used in those experiments. As our model produces both an explanation and final prediction, SFT requires supervision beyond the pairwise assignment labels used in RL. To obtain high-quality supervision, we prompt Gemini-2.5-Flash [gemini2.5] with samples from Twin and retain responses where Gemini is correct. We use default generation parameters for Gemini and collect a total of 136136K high-quality answers.
We train Qwen2.5-VL 3B end-to-end using SFT on the collected samples. We use LLaMa-Factory [llamafactory] for training, using a learning rate of $10^{-6}$ with a cosine scheduler. We train the SFT model for $2$ epochs, as this yielded better accuracy on Fgvqa. All hyperparameters are in Tab. 9.
Figure 10: Examples from Birds-to-Words [birdstowords]. The dataset features 3.33.3K pairs of different bird species, along with a human-annotated description of their differences. The differences are often coarse-grained (e.g. Pair 3, Pair 6 – which both display significant color differences), and the text descriptions subjective (e.g. Pair 4 – where animal 1 is described as ”more beautiful” than animal 2).
@@ Appendix D Failure Cases
Figure 11: Examples from SpotTheDiff [spotthediff]. The dataset features 1313K pairs of images taken from different frames of the same video. Unlike Twin, that emphasizes fine differences in object texture and geometry, the differences in SpotTheDiff are primarily spatio-temporal (e.g. Pair 1 with the blue car leaving the frame, Pair 2 with the people moving) and highly coarse-grained (e.g. Pair 3).
Appendix E Comparisons to Prior Datasets
@@
We provide a more detailed comparison to prior datasets Birds-To-Words [birdstowords] and SpotTheDiff [spotthediff]. While both of these datasets also feature pairs of images as input (similar to Twin), there are several key differences. First, the task in these datasets is to identify the differences between two images, with the assumption a priori that the images are indeed different. In contrast, Twin includes both negative pairs and positive pairs in which the instance is unchanged, requiring models to reason about both similarities and differences rather than only locating discrepancies. Second, the differences captured in Twin are substantially more fine-grained, often involving subtle variations in texture or part geometry that go beyond the relatively coarse differences in prior datasets. Third, our dataset is considerably larger in scale (561561K pairs vs. 1313K for SpotTheDiff and 3.33.3K for Birds-To-Words). We provide examples and describe the differences with each individual dataset in more detail below.
E.1 Birds-To-Words
@@
Birds-to-words [birdstowords] features 3.33.3K pairs of images of different bird species, along with human-annotated descriptions of the differences. We show examples from the dataset in Fig. 10. We observe that, unlike the fine-grained differences in Twin, the differences are often much coarser (e.g. Pair 3 and Pair 6, which both feature significant color differences). Moreover, the human-annotated descriptions can be subjective (e.g. Pair 4 which describes one bird as “more beautiful”).
@@
E.2 Comparison to SpotTheDiff
SpotTheDiff [spotthediff] comprises 1313K pairs of images sourced from two different frames of the same video. The videos used depict large scenes from security camera footage. As a result, the differences between frames are predominantly spatio-temporal, concerning which objects or people are visible or no longer visible in the scene (e.g. a car leaving the frame, a person moving positions). We visualize examples from SpotTheDiff [spotthediff] in Fig. 11. These examples highlight that differences are generally coarse-grained and, unlike the examples in Twin, do not require reasoning about viewpoint variation, part-level geometry, or fine-grained texture changes. Similar to Birds-to-Words, SpotTheDiff assumes the paired images are different, whereas Twin additionally emphasizes recognizing similarities through the inclusion of positive pairs. Finally, Twin is orders of magnitude larger in scale (561K vs. 13K pairs) and introduces challenges not present in SpotTheDiff, including substantial viewpoint variation, lighting changes, and differences in context and background.
@@ E.2 Comparison to SpotTheDiff
Figure 13: Fgvqa Prompts: Ilias [ilias].
Figure 14: Fgvqa Prompts: Landmarks [landmarksv2].
Figure 15: Fgvqa Prompts: Met [met].
Figure 16: Fgvqa Prompts: Cub [cub].
Figure 17: Fgvqa Prompts: Inquire [inquire].

### v2

@@ 5.3 Additional Results
To ensure post-training on Twin does not erode general visual reasoning, we report results on popular VQA benchmarks in Sec. 5.2. These benchmarks span common sense understanding (Seed [39], Mmmu [108], Nlvr2 [78]), spatial reasoning and grounding (Realworld [100], Cv-Bench [83], Blink [19]), hallucination and bias (Pope [41], VlmBias [89]), OCR (TextVQA [77]), and chart/diagram understanding (ChartQA [54], Ai2d [32]). Although not our main focus, models post-trained on Twin retain performance on general VQA benchmarks, with small improvements on most for both Qwen2.5-VL ($1.4\%$ on Nlvr2; $0.8\%$ on Ai2d) and InternVL3.5 ($0.4\%$ on Realworld; $0.4\%$ on Cv-Bench).
Mmlu [23] HellaSwag [111] Gsm8K [12]
Qwen2.5 VL 3B [5] 64.7 66.4 70.8
@@
Beyond vision, Tab. 6 reports performance on popular text-only benchmarks covering different skills: Mmlu [23] for common knowledge, HellaSwag [111] for sentence completion, and Gsm8K [12] for mathematical reasoning. We find that post-training leaves performance largely unchanged on these benchmarks: $64.7\%$ vs $64.6\%$ for Qwen2.5-VL on Mmlu; $57.8\%$ vs $58.1\%$ for InternVL3.5 on Gsm8K. These results reaffirm prior findings that RL-based tuning preserves model skills [75, 11].
@@ A.1 Evaluation Benchmarks
ILIAS [33] is a large-scale test dataset of instance-level image retrieval. It predominantly features images of retail products taken in various contexts, backgrounds, and lighting. For both pair and multi queries, we source images from the core_db split, using the __key__ field to determine instance identity.
Google Landmarks v2 (Landmarks) [97] is a landmark recognition dataset featuring human-made and natural landmarks. The dataset has been used in both classification and retrieval settings. To ensure sufficient images per landmark to for multi queries, we source images from the train split and use the label field to determine landmark identity.
MET [107] is an image retrieval dataset featuring artwork from the Metropolitan Museum of Art in New York. The dataset features images of the same art piece or sculpture from varying viewpoints, emphasizing multi-view consistency in retrieval. We use the mini_met set to construct both pair and multi queries.
CUB [91] is a fine-grained classification dataset that focuses on identifying bird species from images. We use the test split for all queries and create pair and multi queries based on the specie of bird.
Inquire [88] is a benchmark for natural world image retrieval, featuring images of animal and plant species sourced from iNaturalist [87]. We source images from the validation split and use the inat24_species_name field to determine instance identity for both pair and multi queries.
@@ A.2 Fgvqa Prompts
We include all prompts used on all datasets for both pair and multi queries in Figs. 12 to 17. All prompts are identical in format, except for the domain-specific instance names (e.g. bird/object/artwork/species). Additionally, for benchmarks featuring object instances (Twin-Eval, Ilias [33]), we found it necessary to provide the definition of an object instance as models were incorrectly labeling different objects of the same-category (e.g. two earbuds of different colors) as the same.
@@ B.1 Instance Sourcing
We used the Amazon Reviews dataset [24] to extract a large set of product listings, which would serve as candidate objects, and associated noisy review images. We categorized these objects into supercategories based on their associated category in the Amazon store (e.g. “Electronics”) and further into categories based on the product title (e.g. “Speaker”). We aimed to maximize diversity in our instances and thus selected 36 categories to draw from. We subsequently tasked human annotators with removing products that either lacked visual consistency (e.g. the same product being sold in multiple colors) or did not adhere to the object category (e.g. “earbud replacement tips” instead of “earbuds”).
Figure 8: Additional Synthetic Negatives. For each pair, we show a real image of the object instance and a synthetic image of the same instance generated with Dreambooth [70]. We find the synthetic negatives capture the “gist” of the instance, but are not faithful to fine details such as color, geometry, and texture – making them a strong additional set of hard negative pairs.
@@ B.3 Hard Negative Pair Assignment
Once we have collected a set of object instances, we construct hard negative pairs – pairs of distinct objects that appear similar. We collect these hard negatives with the help of human annotators. First, using CLIP [67], for each object instance, we compute pairwise cosine similarities between the image embeddings of that instance and all other instances. We use these similarities to shortlist $k$ visually similar candidates for each instance, with $k$ being a random number between $1-2\times$ the number of positive images of that instance. Human annotators then select the final pairs, omitting pairs they deem to be too easy.
@@ B.4 Additional Details on Generated Negatives
We augment the set of negative pairs with additional synthetically generated images. Augmenting with generated negatives allows us to produce additional negative pairs without extra data collection. We follow the procedure outlined in Dreambooth [70]. For each object instance, we sample 3 to 5 images for each instance to serve as grounding images. We set the initializer tokens to the generic category associated with the instance (e.g. “vase”, “headphones”). Our prompts to Dreambooth are: [“an image of {category}”, “similar to {category}”, “a picture of{category}, “show me a {category}, “here is a {category}”] replacing {category} with the category of the instance. Finally, we task human annotators with validating the generated samples. Annotators are asked to remove generations that either do not appear similar to the instance or are indistinguishable from the instance – yielding a set of hard negative pairs.
@@ Appendix C Additional Training Details
We include additional details for all model training. In Sec. C.1, we provide an overview of GRPO [73], our main post-training method, and our training setup. We provide additional details on supervised fine-tuning (Sec. 5.2) in Sec. C.2. We include all hyperparameters in Tab. 10 and Tab. 11.
@@ C.1 GRPO
To post-train VLMs on Twin, we use Group Relative Policy Optimization (GRPO; [73]). We provide an overview of this method below.
@@
Optimization. We optimize our VLM $\pi_{\theta}$ via GRPO [73]. For each image pair $(I_{1},I_{2})$ with ground truth label $y$, we sample $G$ responses: $\mathcal{O}=\{(I_{1},I_{2},r^{(i)},\hat{y}^{(i)})\}^{G}_{i=1}$ and compute centralized advantages:
@@
Implementation. We train Qwen2.5-VL-3B-Instruct [5] and InternVL3.5-1B-Instruct [94] on Twin, chosen as leading open-source VLMs at the 3B and 1B scales. We do not freeze any part of the model, and train on $4$ A100 GPUs for 1 epoch. We use a batch size of $480$, group size $5$ and learning rate $10^{-6}$. We build on the verl [76] repository for training. We detail all hyperparameters in Tab. 10. Post-training the Qwen2.5-VL 3B model for one epoch on the 560560K samples of TWIN took approximately 140 hours on 4 A100 GPUs.
@@ C.2 Supervised Fine-Tuning
We compare post-training methods in Sec. 5.2. We provide additional details for supervised fine-tuning (SFT) used in those experiments. As our model produces both an explanation and final prediction, SFT requires supervision beyond the pairwise assignment labels used in RL. To obtain high-quality supervision, we prompt Gemini-2.5-Flash [13] with samples from Twin and retain responses where Gemini is correct. We use default generation parameters for Gemini and collect a total of 136136K high-quality answers.
We train Qwen2.5-VL 3B end-to-end using SFT on the collected samples. We use LLaMa-Factory [113] for training, using a learning rate of $10^{-6}$ with a cosine scheduler. We train the SFT model for $2$ epochs, as this yielded better accuracy on Fgvqa. All hyperparameters are in Tab. 11.
Figure 10: Examples from Birds-to-Words [18]. The dataset features 3.33.3K pairs of different bird species, along with a human-annotated description of their differences. The differences are often coarse-grained (e.g. Pair 3, Pair 6 – which both display significant color differences), and the text descriptions subjective (e.g. Pair 4 – where animal 1 is described as ”more beautiful” than animal 2).
@@ Appendix D Failure Cases
InternVL on Inquire. We observe a slight decrease in performance on Inquire ($1.2\%$) when post-training InternVL3.5 1B on Twin. We investigate the model responses and find that post-training the 1B model on Twin reduces “direct identification” (e.g., naming a species) and emphasizes comparisons of visual cues (e.g., the color of the feathers). While this improves instance-level reasoning, it can fail under challenging lighting or occlusion, where the base model succeeds by relying more on priors.
Appendix E Additional Evaluations
We provide additional evaluations for models post-trained with Twin. In Sec. E.1 we explore the impact of post-training on Twin on the related tasks of monument doppelganger detection and evaluate sensitivity to color and shapes. In Sec. E.2 we compare post-training to prompting with in-context-learning examples.
E.1 Additional Benchmarks
We evaluate on an additional set of benchmarks to determine if post-training on TWIN improves robustness to other image variations. We compare the baseline Qwen2.5-VL 3B model with our Twin post-trained variant on monument doppelganger detection [8], the ”yes or no” queries on shape/color sensitivity from VLMs Eye [57], and $500$ pairwise queries from CUTE [34] in Tab. 8.
Doppel. [8] Shape Sens. [57] Color Sens. [57] CUTE [34]
Qwen2.5-VL 3B 50.8 56.8 96.1 58.8
+ Twin 55.2 95.7 99.7 68.0
Table 8: Additional benchmarks.
Post-training on Twin improves doppelganger detection ($4.4\%$), despite Twin not including monuments. This mirrors the gains on Landmarks and Met seen in Sec. 4. Similarly, post-trained models improve at detecting dissimilar shapes ($38.9\%$), reach near-perfect performance at comparing colors ($99.7\%$) and are more robust to photometric variations ($9.2\%$ on CUTE). These results reaffirm that Twin pushes models to attend to subtle visual cues over coarse semantic similarity.
E.2 In-Context Learning
We report results on Fgvqa for in-context-learning (+ICL) prompting in Tab. 9 to explore if we can achieve improved performance via sophisticated prompting as opposed to post-training. We randomly sample one positive and one negative pair from Twin as in-context examples for each Fgvqa query.
Mean TWIN-Eval ILIAS Landmarks MET CUB Inquire
Qwen2.5-VL 3B (+ ICL) 39.5 39.0 37.3 37.5 38.5 43.5 41.2
Qwen2.5-VL 3B (+ TWIN) 67.0 67.3 61.8 57.9 66.0 75.1 73.7
Table 9: In-Context Learning.
Post-training significantly outperforms in-context-learning ($67.0\%$ vs $39.5\%$ on average). In the ICL setting, the 3B model appears to infrequently misunderstand the order of images in the prompt, or ignores the examples entirely.
@@ E.2 In-Context Learning
Figure 11: Examples from SpotTheDiff [29]. The dataset features 1313K pairs of images taken from different frames of the same video. Unlike Twin, that emphasizes fine differences in object texture and geometry, the differences in SpotTheDiff are primarily spatio-temporal (e.g. Pair 1 with the blue car leaving the frame, Pair 2 with the people moving) and highly coarse-grained (e.g. Pair 3).
Appendix F Comparisons to Prior Datasets
@@
We provide a more detailed comparison to prior datasets Birds-To-Words [18] and SpotTheDiff [29]. While both of these datasets also feature pairs of images as input (similar to Twin), there are several key differences. First, the task in these datasets is to identify the differences between two images, with the assumption a priori that the images are indeed different. In contrast, Twin includes both negative pairs and positive pairs in which the instance is unchanged, requiring models to reason about both similarities and differences rather than only locating discrepancies. Second, the differences captured in Twin are substantially more fine-grained, often involving subtle variations in texture or part geometry that go beyond the relatively coarse differences in prior datasets. Third, our dataset is considerably larger in scale (561561K pairs vs. 1313K for SpotTheDiff and 3.33.3K for Birds-To-Words). We provide examples and describe the differences with each individual dataset in more detail below.
F.1 Birds-To-Words
@@
Birds-to-words [18] features 3.33.3K pairs of images of different bird species, along with human-annotated descriptions of the differences. We show examples from the dataset in Fig. 10. We observe that, unlike the fine-grained differences in Twin, the differences are often much coarser (e.g. Pair 3 and Pair 6, which both feature significant color differences). Moreover, the human-annotated descriptions can be subjective (e.g. Pair 4 which describes one bird as “more beautiful”).
@@
F.2 Comparison to SpotTheDiff
SpotTheDiff [29] comprises 1313K pairs of images sourced from two different frames of the same video. The videos used depict large scenes from security camera footage. As a result, the differences between frames are predominantly spatio-temporal, concerning which objects or people are visible or no longer visible in the scene (e.g. a car leaving the frame, a person moving positions). We visualize examples from SpotTheDiff [29] in Fig. 11. These examples highlight that differences are generally coarse-grained and, unlike the examples in Twin, do not require reasoning about viewpoint variation, part-level geometry, or fine-grained texture changes. Similar to Birds-to-Words, SpotTheDiff assumes the paired images are different, whereas Twin additionally emphasizes recognizing similarities through the inclusion of positive pairs. Finally, Twin is orders of magnitude larger in scale (561K vs. 13K pairs) and introduces challenges not present in SpotTheDiff, including substantial viewpoint variation, lighting changes, and differences in context and background.
@@ F.2 Comparison to SpotTheDiff
Figure 13: Fgvqa Prompts: Ilias [33].
Figure 14: Fgvqa Prompts: Landmarks [97].
Figure 15: Fgvqa Prompts: Met [107].
Figure 16: Fgvqa Prompts: Cub [91].
Figure 17: Fgvqa Prompts: Inquire [88].